OpenStudy (anonymous):

A mass of 8.300 kg is suspended from a 1.670 m long string. It revolves in a horizontal circle. The tangential speed of the mass is 2.637 m/s. Calculate the angle between the string and the vertical (in degrees).

OpenStudy (anonymous):

\[T \cos \theta=Mg\] \[T \sin \theta=M \times(V ^{2} \div R)\] \[\tan \theta=V ^{2} \div(R \times G)\] \[\theta=22.6 degrees\]